# Orbital decay candidates reconsidered: WASP-4 b is not decaying and Kepler-1658 b is not a planet

### v2

@@ Abstract
The fate of hot Jupiters is thought to be engulfment by their host stars, the outcome of tidal orbital decay. Transit timing has revealed a few systems with apparently shrinking orbital periods, but such signals can be mimicked by light travel-time effects (LTTE) of a distant companion. Combining transit timings with precise radial-velocity data, including new data, we reassessed three reported cases of orbital decay: WASP-4, WASP-12, and Kepler-1658. For WASP-4, the period change is best explained by LTTE due to a $10~M_{\rm Jup}$ companion at 4 AU, with no need to invoke orbital decay. For WASP-12, in contrast, the data firmly exclude LTTE and confirm genuine orbital decay. For Kepler-1658, spectroscopic and photometric anomalies reveal the “planet” to be an eclipsing K/M binary bound to the F-type primary, with LTTE explaining the observed period change. Thus, among the known hot Jupiters, only WASP-12 b currently shows compelling evidence for orbital decay.
@@ 1 Introduction
Yee et al. (2020) used RV data to rule out LTTE as the cause of WASP-12 b’s shrinking period. Two other systems, WASP-4 b and Kepler-1658 b, also have apparently shrinking periods Although LTTE have been considered in both cases, and some authors have argued that they have been ruled out, no joint analysis of transit times and RVs has yet been presented.
@@ 2 Mathematical preliminaries
$t_{k}=t_{0}+kP_{1}+\frac{z_{2}(t_{0}+kP_{1})-z_{2}(t_{0})}{c},$ (10)
where $t_{0}$ is the transit time of an arbitrarily chosen orbit, $k$ is an integer, and the last term is the LTTE correction.
It will be useful to simplify these equations for the case when the period of the wide-orbiting companion is longer than the time span of the data. We designate $t=0$ as the time of the first data point and perform a Taylor expansion in the small parameter $2\pi t/P_{2}=n_{2}t$:
$\dot{z}_{2}=\sum_{j=0}^{\infty}\beta_{j}(n_{2}t)^{j}.$ (11)
For the special case $e=0$, the first four coefficients are
$\beta_{0}=$ $-K_{2}\sin\phi_{0}$ $~{\rm(velocity),}$ (12)
$\beta_{1}=$ $-K_{2}n_{2}\cos\phi_{0}$ $~{\rm(acceleration),}$ (13)
$\textstyle\beta_{2}=$ $\tfrac{1}{2}K_{2}n_{2}^{2}\sin\phi_{0}$ $~{\rm(jerk),\,and}$ (14)
$\beta_{3}=$ $\tfrac{1}{6}K_{2}n_{2}^{3}\cos\phi_{0}$ $~{\rm(snap),}$ (15)
where $\phi_{0}\equiv\theta(0)-\theta_{\rm c}$. The corresponding approximation for the transit times is obtained using Equation 10 and the time integral of Equation 11:
$t_{k}\approx t_{0}+kP_{1}+\frac{1}{n_{2}c}\sum_{j=1}^{\infty}\frac{k^{j}}{j}\beta_{j-1}(n_{2}P_{1})^{j}.$ (16)
@@ 3.3 Timing+RV analysis
The model had nine free parameters: $K_{1}$, $P_{1}$, $t_{\rm c,1}$, the three $\beta$ parameters, and the three $\gamma$ offsets (one per spectrograph). The best-fit yielded $\chi^{2}=158$ with 164 degrees of freedom, indicating an excellent fit. Posterior sampling was performed using the Markov Chain Monte Carlo (MCMC) algorithm of Goodman & Weare (2010) as implemented in emcee (Foreman-Mackey et al., 2013). To allow for possible underestimation of uncertainties, we introduced a velocity jitter term for each spectrograph. Table 1 gives the results and Figure 1 shows the corresponding fits. The model provides a good fit to all the observables.
@@
$K_{1}$ 237.8 $237.8^{+2.6}_{-2.7}$ m/s
$\beta_{1}$ 16.9-16.9 $16.4^{+1.4}_{-1.4}$ m/s/year
$\beta_{2}$ 4.7-4.7 $4.56^{+0.65}_{-0.66}$ m/s/year2
$\beta_{3}$ $6.5$ $5.3^{+1.6}_{-1.7}$ m/s/year3
$\gamma_{\rm CORALIE}$ 29.1-29.1 $34.0^{+6.4}_{-6.4}$ m/s
$\gamma_{\rm HARPS}$ 60.6-60.6 $62.0^{+7.5}_{-7.9}$ m/s
$\gamma_{\rm HIRES}$ 44.9-44.9 $44.4^{+3.8}_{-3.8}$ m/s
$\sigma_{\rm CORALIE}$ $6.5$ $9.4^{+5.7}_{-5.9}$ m/s
$\sigma_{\rm HARPS}$ $7.3$ $10.0^{+5.2}_{-3.9}$ m/s
$\sigma_{\rm HIRES}$ $2.2$ $5.4^{+3.0}_{-2.1}$ m/s
$K_{2}$ $23.0$ $22.5^{+7.9}_{-2.4}$ m/s
$P_{2}$ $7.2$ $7.10^{+2.03}_{-0.74}$ years
$a_{2}$ $3.6$ $3.56^{+0.60}_{-0.23}$ AU
$m_{2}\sin I_{2}$ $9.4$ $9.2^{+3.8}_{-1.2}$ $M_{\rm Jup}$
Note. — Parameters below the line were constructed from parameters above the line assuming $e=0$. Data in Column 3 are based on the 16th, 50th, and 84th percentiles of the cumulative posterior probability function for each parameter, marginalized over all other parameters.
@@
$n_{2}=\sqrt{-\frac{2\beta_{3}}{\beta_{1}}}~~{\rm and}~~K_{2}=\frac{\beta_{1}}{n_{2}}\sqrt{1+\left(\frac{2\beta_{2}}{n_{2}^{2}}\right)^{\!\!2}}.$ (19)
Adopting a stellar mass of $0.9\,M_{\odot}$ (Wilson et al., 2008), we infer a companion with minimum mass $\sim$10 $M_{\rm Jup}$ at $\sim$4 AU (Table 1, Figure 2). The posterior distributions for $a_{2}$ and $m_{2}\sin i_{2}$ have long tails extending toward larger masses and wider separations, but the radial acceleration of the primary star is constrained more tightly:
$\frac{m_{2}\sin I_{2}}{a_{2}^{2}}=0.704_{-0.029}^{+0.068}~\frac{M_{\rm Jup}}{{\rm AU}^{2}}.$ (20)
@@
In summary, a super-Jupiter or brown dwarf companion at several AU naturally explains the transit-timing and RV variations of WASP-4, without invoking orbital decay.
@@ 5 Kepler-1658
Kepler-1658 was first described in detail by Chontos et al. (2019), using data from NASA’s Kepler mission (Borucki et al., 2010) and follow-up observations. Early in the mission, a sequence of transits was detected with a period of 3.9 days and a depth of 0.13%. The source was inconsistently classified in mission catalogs, appearing as a planet candidate in some and a false positive in others. Confirmation of Kepler-1658 b as a planet was based mainly on the detection of radial-velocity variations with the expected period and phase and an amplitude consistent with a planetary mass ($\approx$ 6 $M_{\rm Jup}$). The host star is a subgiant with effective temperature 6200 K, mass $1.4\,M_{\odot}$, and radius $2.9\,R_{\odot}$. Because tidal dissipation rates are expected to increase as a star evolves into the subgiant phase, the system was highlighted as a favorable target for detecting orbital decay.a
@@ 7 Acknowledgments
J.N.W. gratefully acknowledges support from the NASA TESS project and a NASA Keck PI award. The work described in this paper was based in part on observations with the NEID instrument on the WIYN 3.5m telescope at Kitt Peak National Observatory. Kitt Peak is a facility of NSF’s NOIRLab, managed by the Association of Universities for Research in Astronomy (AURA). The WIYN telescope is a joint facility of NOIRlab, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, Purdue University, and Princeton University. NEID was funded by the NASA-NSF Exoplanet Observational Research (NN-EXPLORE) partnership and built by Pennsylvania State University. The NEID archive is operated by the NASA Exoplanet Science Institute at the California Institute of Technology. NN-EXPLORE is managed by the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration. The authors thank the NEID Queue Observers and WIYN Observing Associates, as well as the California Planet Search team supporting precise RV acquisition with the Keck I telescope, for their dedicated and skillful execution of the observations.

### v3

@@ Abstract
The fate of hot Jupiters is thought to be engulfment by their host stars, the outcome of tidal orbital decay. Transit timing has revealed a few systems with apparently shrinking orbital periods, but such signals can be mimicked by light travel-time effects (LTTE) of a distant companion. Combining transit timings with precise radial-velocity data, including new data, we reassessed three reported cases of orbital decay: WASP-4, WASP-12, and Kepler-1658. For WASP-4, the period change is best explained by LTTE due to an $\approx$$8~M_{\rm Jup}$ companion at $\approx$8 AU, with no need to invoke orbital decay. For WASP-12, in contrast, the data firmly exclude LTTE and confirm genuine orbital decay. For Kepler-1658, spectroscopic and photometric anomalies reveal the “planet” to be an eclipsing K/M binary bound to the F-type primary, with LTTE explaining the observed period change. Thus, among the known hot Jupiters, only WASP-12 b currently shows compelling evidence for orbital decay.
@@ 1 Introduction
Yee et al. (2020) used RV data to rule out LTTE as the cause of WASP-12 b’s shrinking period. Two other systems, WASP-4 b and Kepler-1658 b, also have apparently shrinking periods. Although light travel-time effects have been considered in both cases, and some authors have argued that they have been ruled out, no joint analysis of transit times and RVs has yet been presented.
@@ 2 Mathematical preliminaries
$t_{k}=t_{\rm c,1}+kP_{1}+\frac{z_{2}(t_{\rm c,1}+kP_{1})-z_{2}(t_{\rm c,1})}{c},$ (10)
where $t_{\rm c,1}$ is the time of conjunction of an arbitrarily chosen orbit, $k$ is an integer, and the last term is the LTTE correction.
It will be useful to simplify these equations for the case when the period of the wide-orbiting companion is longer than the time span of the data. We choose $t=0$ to be within the time span of the data, and perform a Taylor expansion in the small parameter $2\pi t/P_{2}=n_{2}t$:
$\dot{z}_{2}=\sum_{j=0}^{\infty}\beta_{j}\,t^{j}=\sum_{j=0}^{\infty}\frac{\beta_{j}}{n_{2}^{j}}\,(n_{2}t)^{j}$ (11)
For the special case $e_{2}=0$, the first four coefficients are
$\beta_{0}=K_{2}\sin\phi_{0},$ $~~\beta_{1}=-K_{2}n_{2}\cos\phi_{0}$ (12)
$\beta_{2}=-\tfrac{1}{2}K_{2}n_{2}^{2}\sin\phi_{0},$ $~~{\rm and}~~\beta_{3}=\tfrac{1}{6}K_{2}n_{2}^{3}\cos\phi_{0},$ (13)
where $\phi_{0}\equiv n_{2}t_{\rm c,2}$. The corresponding approximation for the transit times is obtained using Equation 10 and the time integral of Equation 11:
$t_{k}\approx t_{0}+k\left(\!P+\frac{\beta_{0}}{c}\!\right)+\frac{1}{c}\sum_{j=1}^{\infty}\beta_{j}\frac{(t_{0}+kP)^{j+1}-t_{0}^{j+1}}{j+1},$ (14)
and the $\beta_{0}$ term is usually negligible.
@@ 3.3 Timing+RV analysis
The model had nine free parameters: $K_{1}$, $P_{1}$, $t_{\rm c,1}$, the three $\beta$ parameters, and the three $\gamma$ offsets (one per spectrograph). The best-fit yielded $\chi^{2}=184$ with 164 degrees of freedom, indicating an acceptable fit. Posterior sampling was performed using the Markov Chain Monte Carlo (MCMC) algorithm of Goodman & Weare (2010) as implemented in emcee (Foreman-Mackey et al., 2013). To allow for possible underestimation of uncertainties, we introduced a velocity jitter term for each spectrograph. Table 1 gives the results and Figure 1 shows the corresponding fits. The model provides a good fit to all the observables.
@@
$K_{1}$ $237.6$ $237.6^{+2.7}_{-2.7}$ m/s
$\beta_{1}$ 17.1-17.1 $17.1^{+1.2}_{-1.3}$ m/s/yr
$\beta_{2}$ 1.05-1.05 $1.06^{+0.17}_{-0.16}$ m/s/yr2
$\beta_{3}$ $0.171$ $0.174^{+0.041}_{-0.041}$ m/s/yr3
$\gamma_{\rm CORALIE}$ 35.2-35.2 $35.1^{+6.8}_{-6.8}$ m/s
$\gamma_{\rm HARPS}$ 63.8-63.8 $63.3^{+8.5}_{-9.2}$ m/s
$\gamma_{\rm HIRES}$ 46.2-46.2 $46.1^{+3.5}_{-3.7}$ m/s
$\sigma_{\rm CORALIE}$ $10.1$ $10.1^{+6.7}_{-6.3}$ m/s
$\sigma_{\rm HARPS}$ $11.5$ $11.5^{+9.6}_{-4.9}$ m/s
$\sigma_{\rm HIRES}$ $5.3$ $5.3^{+2.9}_{-2.0}$ m/s
$a_{2}$ $8.39$ $7.95^{+0.98}_{-0.41}$ AU
$P_{2}$ $25.6$ $23.6^{+4.5}_{-1.9}$ yr
$K_{2}$ $78.1$ $73.6^{+12.2}_{-4.7}$ m/s
$m_{2}\sin I_{2}$ $7.56$ $6.94^{+1.59}_{-0.60}$ $M_{\rm Jup}$
Note. — Parameters below the line were constructed from parameters above the line assuming $e_{2}=0$. Data in Column 3 are based on the 16th, 50th, and 84th percentiles of the cumulative posterior probability function for each parameter, marginalized over all other parameters.
@@
$n_{2}=\sqrt{-\frac{6\beta_{3}}{\beta_{1}}}~~{\rm and}~~K_{2}=\frac{1}{n_{2}}\sqrt{\beta_{1}^{2}+\frac{4\beta_{2}^{2}}{n_{2}^{2}}}.$ (17)
Adopting a stellar mass of $0.9\,M_{\odot}$ (Wilson et al., 2008), we infer a companion with minimum mass of 6–9 $M_{\rm Jup}$ at 7.5–9 AU (Table 1, Figure 2). The posterior distributions for $a_{2}$ and $m_{2}\sin I_{2}$ have long tails extending toward larger masses and wider separations, but the radial acceleration of the primary star is constrained more tightly:
$\frac{m_{2}\sin I_{2}}{a_{2}^{2}}=0.1041^{+0.0086}_{-0.0042}\frac{M_{\rm Jup}}{{\rm AU}^{2}}.$ (18)
@@
In summary, a super-Jupiter or brown dwarf companion with an orbital separation of about 8 AU and a period of about 25 years naturally explains the transit-timing and RV variations of WASP-4, without invoking orbital decay.
@@ 5 Kepler-1658
Kepler-1658 was first described in detail by Chontos et al. (2019), using data from NASA’s Kepler mission (Borucki et al., 2010) and follow-up observations. Early in the mission, a sequence of transits was detected with a period of 3.9 days and a depth of 0.13%. The source was inconsistently classified in mission catalogs, appearing as a planet candidate in some and a false positive in others. Confirmation of Kepler-1658 b as a planet was based mainly on the detection of radial-velocity variations with the expected period and phase, an amplitude consistent with a planetary mass ($\approx$ 6 $M_{\rm Jup}$), and no evidence for line-bisector variations. The host star is a subgiant with effective temperature 6200 K, mass $1.4\,M_{\odot}$, and radius $2.9\,R_{\odot}$. Because tidal dissipation rates are expected to increase as a star evolves into the subgiant phase, the system was highlighted as a favorable target for detecting orbital decay.
@@ 7 Acknowledgments
We thank Shreyas Vissapragada, Dan Huber, and Özgür Baştürk for feedback on the manuscript. J.N.W. gratefully acknowledges support from the NASA TESS project and a NASA Keck PI award. The work described in this paper was based in part on observations with the NEID instrument on the WIYN 3.5m telescope at Kitt Peak National Observatory. Kitt Peak is a facility of NSF’s NOIRLab, managed by the Association of Universities for Research in Astronomy (AURA). The WIYN telescope is a joint facility of NOIRlab, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, Purdue University, and Princeton University. NEID was funded by the NASA-NSF Exoplanet Observational Research (NN-EXPLORE) partnership and built by Pennsylvania State University. The NEID archive is operated by the NASA Exoplanet Science Institute at the California Institute of Technology. NN-EXPLORE is managed by the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration. The authors thank the NEID Queue Observers and WIYN Observing Associates, as well as the California Planet Search team supporting precise RV acquisition with the Keck I telescope, for their dedicated and skillful execution of the observations.